# Out-of-Distribution Detection using Neural Activation Prior

### 01 Mar 2024 v2

@@ Abstract
Out-of-distribution detection is a crucial technique for deploying machine learning models in the real world to handle the unseen scenarios. In this paper, we propose a simple but effective Neural Activation Prior (NAP) for out-of-distribution detection (OOD). Our neural activation prior is based on a key observation that, for a channel before the global pooling layer of a fully trained neural network, the probability of a few of its neurons being activated with a larger response by an in-distribution (ID) sample is significantly higher than that by an OOD sample. An intuitive explanation is each channel in a model fully trained on ID dataset would play a role in detecting a certain pattern in the samples within the ID dataset, and a few neurons can be activated with a large response when the pattern is detected in an input sample. Thus, a new scoring function based on this prior is proposed to highlight the role of these strongly activated neurons in OOD detection. This approach is plug-and-play and does not lead to any performance degradation on in-distribution data classification and requires no extra training or statistics from training or external datasets. Notice that previous methods primarily rely on post-global-pooling features of the neural networks, while the within-channel distribution information we leverage would be discarded by the global pooling operator. Consequently, our method is orthogonal to existing approaches and can be effectively combined with them in various applications. Experimental results show that our method achieves the state-of-the-art performance on CIFAR-10, CIFAR-100 and ImageNet datasets, which demonstrates the power of the proposed prior.
@@ 1 Introduction
Recently, deep learning has developed rapidly and become a crucial technique in various fields. However, neral networks frequently encounter data in inference that differs greatly from their training data, which is known as out-of-distribution (OOD) data. This challenge is growing more prevalent and is particularly vital in safety-critical areas such as autonomous driving [9, 16] and medical diagnosis [29], which urges the development of effective OOD detection methods.
@@
Figure 1: Activation distribution difference between ID data and OOD data inside channel. This illustrates the difference in within-channel activation distribution between CIFAR-10 [19] (ID data) and the Texture [5] (OOD data) datasets. The horizontal axis represents the average activation value of the channel, while the vertical axis shows the maximal activation value of the channel. Additionally, semi-transparent shaded areas are introduced in the graph to represent the uncertainty surrounding each data point. As depicted, in this two-dimensional space formed by mean and maximum values, the ID and OOD data are linearly separable. Moreover, it is observed that, for identical mean values, the in-distribution data exhibit notably higher maximal activation values within the channel compared to the out-of-distribution data.
@@
Figure 2: Visualization of different score distributions obtained using Densenet [13] on ID dataset (CIFAR-10 [19]) and OOD datasets (iSun [40] and SVHN [27]) By simply multiplying (a) Energy Score with (b) NAP Score, the (c) Energy $\times$ NAP Score can achieve better results.This is because our proposed Neural Activation Prior is orthogonal to a series of existing OOD detection methods represented by Energy Score, so better results can be obtained by simple multiplication. It should be noted that our intention is not to compete with the Energy Score, but rather to emphasize that our proposed scoring function is complementary to it, as well as to other methods based on the Energy Score.
In the domain of out-of-distribution detection, OOD data refers to types of data that the model has not encountered during training. In practice, OOD data always has large diversity and is difficult to identify [41]. Researchers often phrase OOD detection as a one-class classification task, utilizing prior knowledge. Such methods [11, 23, 43] mainly rely on various priors to construct scoring functions to distinguish OOD samples from in-distribution samples. For example, Hendrycks et al. [11] observed that out-of-distribution samples typically exhibit lower maximum softmax probabilities, and accordingly proposed using the maximum softmax probability output by a neural network as an OOD indicator. Liu et al. [23] found that out-of-distribution samples usually have lower logits values, and based on this, an energy function was proposed for OOD detection. Although these methods have shown some promising results, there is still a gap in their application in real-world scenarios. At present, a mature and definitive paradigm for OOD detection has not yet established, encouraging us to explore various aspects of the OOD detection problem from different perspectives. Thus, in this paper, we aim to explore an effective OOD detection method to provide new insights into the field.
In this study, we propose a novel prior, called Neural Activation Prior (NAP), for out-of-distribution detection. Our neural activation prior is based on a key observation that for channels before the global pooling layer in a fully trained neural network (illustrated to Figure 3), a few neurons are significantly more likely to show a larger response when activated by in-distribution (ID) samples compared to OOD samples. An intuitive explanation is that channels in a model fully trained on an ID dataset play a role in detecting certain patterns in the input samples from the ID dataset. When such patterns are detected, a few neurons can be activated, resulting in larger responses. These large responses usually occurs when the input sample is ID data, but when the input is OOD data, such responses are rarely observed. This is because the pattern that the neuron is focusing on is unlikely to be present in the OOD data. To verify our proposed prior, we conduct experiments using the Densenet architecture [13] on the CIFAR-10 [19] and Texture [5] datasets. We compute the mean and maximum values of the within-channel activations before global pooling of all samples at the penultimate layer. As shown in Figure 1, it is obvious that the ID samples and OOD samples are linearly separable in the figure, and when the average activation value is the same, the maximal activation value of ID samples are significantly larger than that of OOD data.
It is worth noting that our proposed prior is orthogonal to that used in current OOD detection methods. In the field of out-of-distribution detection, as shown in Figure 3, existing methods [11, 22, 23, 34, 33, 8, 2] mainly focus on the outputs and weights of the neural network after global pooling, and use them to design scoring functions for OOD detection. In contrast, the prior we proposed is focus within the channels of the penultimate layer before global pooling. Since the information carried in our prior can be easily lost during the global pooling process, this demonstrates that our NAP is essentially complementary to the priors used in current OOD detection studies. This property is verified in subsequent experiments Section 4 in this paper.
Furthermore, in this paper, we propose a simple yet effective scoring function based on our prior NAP for OOD detection. To be precise, our scoring function is based on the ratio of the maximal and averaged activation values within a channel. It is a natural definition. On one hand, inspired by the concept of Signal-to-Noise Ratio (SNR), we can consider the maximal activation value as signal strength, while the averaged activation value represents noise strength. Therefore, the ratio of the maximum to the average value can be used to measure the quality of information contained in the channel. On the other hand, as shown in Figure 1, the ratio of the maximal and the averaged values for in-distribution samples is significantly higher than that of out-of-distribution data. Hence, from both the physical and practical standpoints, using this ratio as an OOD scoring function is a reasonable choice. It is noteworthy that this scoring function is a plug-and-play method, requiring no additional training, extra data, or reliance on pre-calculated statistical data from the training set. It does not lead to any degradation of the model’s inherent classification capabilities, and is complementary to existing methods and can be effectively combined with them to obtain better OOD detection results. For example, we can simply multiply our proposed scoring function with the energy function, and then we can significantly improve the OOD detection effect. Figure 2 shows a visualization of this result. These attributes greatly enhance the practicality of our approach, making our OOD detection technique applicable in a wide range of scenarios.
@@
Experimental results show that our method achieves state-of-the-art performance on CIFAR-10 [19], CIFAR-100 [19] and ImageNet datasets. Specifically, our method significantly reduces the false positive rate by 40.06% (from 15.05 to 9.02) on the CIFAR-10 dataset [19], a reduction of 37.89% (from 41.40 to 25.71) on the CIFAR-100 dataset [19], and a reduction of 11.69% (from 35.66 to 31.49) on the ImageNet dataset. The large drop in FPR95 highlights the effectiveness of our approach in different environments. The above experimental results demonstrate the power of our proposed prior. We believe that these findings will provide inspiration for other researchers and thus promote progress in the field of OOD detection.
@@
  • We propose a novel prior in OOD detection, named Neural Activation Prior (NAP). NAP is orthogonal to that used in existing OOD detection methods and we believe this prior can inspire researchers to further explore better OOD detection methods.
  • Based on the proposed prior, we develop a simple yet effective OOD detection scoring function. It is a plug-and-play approach that does not impair the model’s inherent capabilities. It can be readily integrated with many existing OOD detection techniques, enhancing their ability to balance OOD detection with in-distribution accuracy.
  • We demonstrate the state-of-the-art performance of our approach through extensive experiments across various datasets, highlighting its efficiency, ease of implementation, and effectiveness.
Discussion.
  • The scoring function proposed in this paper is a preliminary attempt to utilize the prior we proposed, aimed primarily at demonstrating its potential in OOD detection. There might be more effective scoring functions derived from this prior, but that is not the main concern of this paper.
  • We emphasize that the scoring function proposed in this paper is not intended to compete with functions proposed in existing work. Based on our proposed prior NAP, this scoring function is complementary to various existing works (such as a series of works [23, 34, 8] based on Energy Score). Therefore, our main focus is on integrating our approach with existing methods.
@@ 2.1 Out-of-distribution detection
Research on out-of-distribution detection has explored a variety of techniques to underscore the distinctions between ID and OOD samples. These methods encompass classification-based [15, 22, 3, 7, 10], density-based [47, 1, 26, 46, 17, 28, 18, 31], and distance-based approaches [35, 21, 25, 30, 4, 14, 44, 37, 36, 24], with classification-based techniques generally outperforming the other types [41]. In classification-based methods, the basic work of OOD detection starts with a simple and effective baseline: using the Maximum Softmax Probability (MSP) [10] to measure the probability that a certain sample is an in-distribution sample. On this basis, early approaches [22, 12, 23] focused on developing enhanced OOD indicators derived from neural network outputs. In addition, some researchers have proposed strategies involving OOD sample generation [20] and gradient-based [22] techniques. Among these, certain post hoc methods [11, 22, 23, 34, 33, 8, 43, 2] are notable for their simplicity and because they do not necessitate changes in the training process or objectives. This feature is particularly valuable for implementing OOD detection in real production environments, where the additional cost and complexity associated with retraining would be unacceptable.
@@
The Maximum Softmax Probability (MSP) method, initially presented by Hendrycks et al. [11], was a formative step in post hoc OOD detection, using a neural network’s softmax output as a heuristic for distinguishing ID from OOD samples. Its straightforward application facilitated early adoption in OOD studies. Despite MSP’s influence, its limitations prompted further innovation, giving rise to the Energy method. This method, proposed by Liu et al. [23], refines the approach by assigning an energy score to network outputs, showing quantitative improvements over MSP with theoretical and empirical support. Advancements in post hoc OOD detection have led to diverse methodological branches stemming from MSP [10] and Energy [23] paradigms. LINe [2] innovates by reducing neuron-induced noise through the calculation of Shapley values. Yu et al. [43] distinguish ID from OOD data by identifying neural network blocks with optimal differentiation based on the norms of their features. DICE [33] improves discrimination by pruning weights in the fully connected layer according to the contribution units make during classification. On the other end of the spectrum, entirely computation-free post hoc methods such as ReAct [34] and ASH [8] have shown promise. ReAct [34] investigates activations prior to the fully connected layer, applying rectification to suppress extreme activations that OOD data tend to trigger, thereby achieving refined detection outcomes. Similarly, ASH [8] prunes the activations inputted to the fully connected layer, but it achieves even more enhanced results compared to DICE [33] by its selective pruning strategy. In this paper, our comparison mainly focuses on post hoc methods, since our method also belongs to this category.
@@ 3 Neural Activation Prior
The neural activation prior is based on the following observation for OOD detection: for a channel located before the global pooling layer in a fully trained neural network, the likelihood that a small number of its neurons activated with a stronger response to an in-distribution (ID) sample is significantly higher compared to an out-of-distribution (OOD) sample. For the behavior in other layers of the neural network, refer to the discussion around Figure 5 in Section 5.5.2.
To formally describe this observation, we first define the concept of neural activation. Consider a trained classification neural network $f$, assuming it receives $D$-dimensional input $x$ and outputs $K$-dimensional logits. That is $f:\mathbb{R}^{D}\rightarrow\mathbb{R}^{K}$. We concentrate on the activation tensor $\mathbf{A}$ , located at the penultimate layer just before the global pooling operation, as illustrated in Figure 3. Let the dimensions of $\mathbf{A}$ be $C\times H\times W$, where $C$ is the number of channels, and $H$ and $W$ are the spatial dimensions.
Figure 3: Illustration identifying the focus zone of Neural Activation Prior (NAP) in a classification neural network. The figure highlights the specific position of NAP in the network. The activation value in the green part is the location of the NAP proposed in this article, which is before global pooling. In contrast, most existing methods primarily focus on the regions after global pooling.
@@
    where $\text{Max}(\mathbf{A}_{j})$ is the maximum value among all elements in the activation vector $\mathbf{A}_{j}$. Notice that $\mathbf{A}_{jkl}$ is processed by ReLU (as shown in Figure 3), hence $\mathbf{A}_{jkl}\geq 0$.
@@
Inspired by the concept of signal-to-noise ratio, we interpret $\text{Max}(\mathbf{A}_{j})$ as the signal strength and $\text{Mean}(\mathbf{A}_{j})$ as the noise strength. The ratio of these values can be viewed as an indicator of the quality of the activation vector $\mathbf{A}_{j}$. Note that we calculate this metric separately for each channel in the network, since different channels are usually used to detect different objects. In the context of OOD detection, this quality measure can be used to assess whether the neural network recognizes the input sample — in other words, this measure can be used to judge whether a sample is within the training set distribution. As shown in Figure 1, this phenomenon is termed as the neural activation prior.
It is worth noting that our proposed prior is orthogonal to existing OOD detection methods. As illustrated in Figure 3, existing methods mainly focus on the network output and weight of the penultimate layer after global pooling, and leverage them to design various scoring functions for OOD detection. In contrast, the prior we propose focuses on the channels of the penultimate layer before global pooling. Since the distribution information within these channels is inevitably lost during the global pooling process, our proposed prior is complementary to existing work.

### 06 Mar 2024 v3

@@ Abstract
Out-of-distribution detection (OOD) is a crucial technique for deploying machine learning models in the real world to handle the unseen scenarios. In this paper, we first propose a simple yet effective Neural Activation Prior (NAP) for OOD detection. Our neural activation prior is based on a key observation that, for a channel before the global pooling layer of a fully trained neural network, the probability of a few neurons being activated with a large response by an in-distribution (ID) sample is significantly higher than that by an OOD sample. An intuitive explanation is that for a model fully trained on ID dataset, each channel would play a role in detecting a certain pattern in the ID dataset, and a few neurons can be activated with a large response when the pattern is detected in an input sample. Then, a new scoring function based on this prior is proposed to highlight the role of these strongly activated neurons in OOD detection. Our approach is plug-and-play and does not lead to any performance degradation on ID data classification and requires no extra training or statistics from training or external datasets. Notice that previous methods primarily rely on post-global-pooling features of the neural networks, while the within-channel distribution information we leverage would be discarded by the global pooling operator. Consequently, our method is orthogonal to existing approaches and can be effectively combined with them in various applications. Experimental results show that our method achieves the state-of-the-art performance on CIFAR benchmark and ImageNet dataset, which demonstrates the power of the proposed prior. Finally, we extend our method to Transformers and the experimental findings indicate that NAP can also significantly enhance the performance of OOD detection on Transformers, thereby demonstrating the broad applicability of this prior knowledge.
Keywords:
Out-of-Distribution Detection Global Pooling Layer Feature Activation Analysis
@@ 1 Introduction
Deep learning has developed rapidly in the last decade and become a crucial technique in various fields. However, neural networks would frequently make erroneous judgments in inference when encounter the data that differs greatly from their training data, which is known as out-of-distribution (OOD) data. This challenge is growing more prevalent and is particularly vital in safety-critical areas such as autonomous driving [10, 17] and medical diagnosis [30], which urges the development of effective OOD detection methods.
In practice, OOD data exhibits large diversity and is difficult to identify [42]. Existing studies typically formulate OOD detection as a one-class classification task, utilizing prior knowledge. They [11, 24, 44] propose various priors, based on which they further design scoring functions to distinguish OOD samples from ID samples. For example, Hendrycks et al. [11] observed that OOD samples always exhibit lower maximum softmax probabilities, and accordingly proposed using the maximal softmax probability output by a neural network as an OOD indicator. Liu et al. [24] found that OOD samples usually have lower logits values, and based on this, an energy function was proposed for OOD detection. Drawing from these precedents, it’s clear that existing methods largely rely on the introduction of certain priors. While some promising results highlight the effectiveness of these heuristics, a gap remains in meeting the practical requirements of real-world applications. Importantly, we note that the focus of these priors is narrowly concentrated on features and weights following global pooling, while the characteristics before the pooling layer are consistently ignored. Thus, we believe that finding and incorporating priors that can complement these existing focuses is essential, which will constitute the main contribution of our work.
In this study, we propose a novel prior, called Neural Activation Prior (NAP), for OOD detection. NAP characterizes our key observation that for channels before the global pooling layer in a fully trained neural network (illustrated to Figure 2(a)), a few neurons have a significantly high probability to show a larger response when activated by in-distribution (ID) samples compared to OOD samples. An intuitive explanation for NAP is that channels in a model fully trained on an ID dataset play a role in detecting certain patterns in the input samples from the ID dataset. When such patterns are detected, a few neurons can be activated [13], resulting in larger responses. These large responses usually occur when the input sample is ID data, but when the input is OOD data, such responses are rarely observed. This is because the pattern that the neuron focuses on is unlikely to be present in the OOD data. To verify our proposed prior, we employed DenseNet architecture [15] on CIFAR-10 [20] and Texture [5] datasets, analyzing mean and maximum within-channel activations before global pooling at the penultimate layer. Figure 1 clearly demonstrates that ID samples showing significantly higher maximal activation values than OOD samples at equal average activation levels. This is not unique to the $333^{\text{rd}}$ channel but is observed across most channels.
@@
Figure 1: Activation distribution difference between ID data and OOD data inside channel. This illustrates the difference in within-channel activation distribution between CIFAR-10 [20] (ID data) and the Texture [5] (OOD data) datasets. The horizontal axis represents the average activation value of the channel, while the vertical axis shows the maximal activation value of the channel. Interestingly, at the same mean activation levels, the ID data (CIFAR-10) show significantly higher maximum activation values within the channel compared to the OOD data (Texture). It’s important to note that this pattern is not unique to the $333^{\text{rd}}$ channel but is observed across most channels.
It is worth noting that our proposed prior is orthogonal to that used in current OOD detection methods. In the OOD deteciton field, as shown in Figure 2(a), existing methods [11, 23, 24, 33, 34, 8] mainly focus on the outputs and weights of the neural network after global pooling, and use them to design scoring functions for OOD detection. In contrast, the prior we proposed is focus within the channels of the penultimate layer before global pooling. Since the information carried in our prior can be easily lost during the global pooling process, this demonstrates that our NAP is essentially complementary to the priors used in current OOD detection studies. This property is verified in subsequent experiments Section 6 in this paper. To this end, we would like to emphasize that the contribution of this paper would lie in how much improvement we can achieve by integrating our method with existing approaches, rather than a direct comparison with existing methods in the field.
Furthermore, in this paper, we propose a simple yet effective scoring function for OOD detection based on our prior NAP. To be precise, our scoring function is based on the ratio of the maximal and averaged activation values within a channel. The rationale behind the scoring function can be understood from two primary perspectives: conceptual inspiration from the Signal-to-Noise (SNR) and empirical validation. On the conceptual front, inspired by the concept of SNR, we can consider the maximal activation value as signal strength, while the averaged activation value represents noise strength. Therefore, the ratio of the maximum to the average value can be used to measure the quality of information contained in the channel. On the empirical front, as shown in Figure 1, the ratio of the maximal and the averaged values for ID samples is significantly higher than that of OOD data. Regarding practical deployment, as previously mentioned, our scoring function complements and, when multiplied with existing metrics, improves OOD detection, as depicted in Figure 2. Also, it is noteworthy that this scoring function is a plug-and-play method, requiring no additional training, extra data, or reliance on pre-calculated statistical data from the training set, which makes it broadly applicable.
@@
Figure 2: Score Distribution Visualization Using DenseNet on CIFAR-10 (ID) and iSun (OOD) The integration of (a) Energy Score and (b) NAP Score through multiplication yields the (c) Energy $\times$ NAP Score, demonstrating superior differentiation between ID and OOD datasets. The effectiveness of this approach is attributed to the orthogonal nature of the proposed NAP relative to conventional OOD detection methods exemplified by the Energy Score. This illustrates that a simple multiplicative combination with NAP enhances detection capability. Importantly, the objective here is not to outperform the Energy Score per se but to highlight the synergistic potential of NAP as a complementary enhancement to the Energy Score and similar methods.
@@
Experimental results show that our method achieves state-of-the-art performance on CIFAR-10 [20], CIFAR-100 [20] and ImageNet datasets. Specifically, our method significantly reduces the false positive rate by 40.06% (from 15.05 to 9.02) on the CIFAR-10 dataset [20], a reduction of 37.89% (from 41.40 to 25.71) on the CIFAR-100 dataset [20], and a reduction of 16.26% (from 35.66 to 29.86) on the ImageNet dataset. The large drop in FPR95 highlights the effectiveness of our approach in different environments. The above experimental results demonstrate the power of our proposed prior. We believe that these findings will provide inspiration for other researchers and thus promote progress in the field of OOD detection.
Additionally, while convolutional neural networks (CNNs) are predominantly employed in OOD detection tasks, the advent of Transformer architectures [39] and their variants has showcased substantial efficacy across a diverse array of applications. Motivated by this, we extend our method to ensure compatibility with Transformer models. The empirical evidence obtained from our experiments affirms the robustness of our approach, demonstrating its adaptability to various architectural paradigms.
@@
  • We introduce the Neural Activation Prior (NAP), a novel contribution to OOD detection. Uniquely, NAP is orthogonal to priors utilized in existing methods, offering a distinct and complementary perspective that paves the way for advanced OOD detection research.
  • Based on the proposed prior, we develop a simple yet effective OOD detection scoring function. It can be readily integrated with many existing OOD detection techniques, enhancing their ability to balance OOD detection with ID accuracy.
  • We demonstrate the state-of-the-art performance of our approach through extensive experiments across various datasets, including a reduction in FPR95 by up to 40.06%. These results underscore the method’s operational efficiency, simplicity of deployment, and overall efficacy.
  • We extend the method to accommodate Transformer architectures. Experimental results are encouraging, validating the method’s efficacy across various architectural designs.
@@ 2.1 OOD Detection
The OOD detection community has explored a variety of techniques to underscore the distinctions between ID and OOD samples. These methods encompass classification-based [16, 23, 3, 7, 12], density-based [48, 1, 27, 47, 18, 29, 19, 31], and distance-based approaches [35, 22, 26, 4, 45, 37, 36, 25], with classification-based techniques generally outperforming the other types [42]. In classification-based methods, the basic work of OOD detection starts with a simple and effective baseline: using the Maximum Softmax Probability (MSP) [12] to measure the probability that a certain sample is an ID sample. On this basis, early approaches [23, 14, 24] focused on developing enhanced OOD indicators derived from neural network outputs. In addition, some researchers have proposed strategies involving OOD sample generation [21] and gradient-based [23] techniques. Among these, certain post-hoc methods [11, 23, 24, 33, 34, 8, 44] are notable for their simplicity and because they do not necessitate changes in the training process or objectives. This feature is particularly valuable for implementing OOD detection in real production environments, where the additional cost and complexity associated with retraining would be unacceptable.
@@
The MSP method, initially presented by Hendrycks et al. [11], was a formative step in post hoc OOD detection, using a neural network’s softmax output as a heuristic for distinguishing ID from OOD samples. Its straightforward application facilitated early adoption in OOD studies. Despite MSP’s influence, its limitations prompted further innovation, giving rise to the Energy method. This method, proposed by Liu et al. [24], refines the approach by assigning an energy score to network outputs, showing quantitative improvements over MSP with theoretical and empirical support. Advancements in post hoc OOD detection have led to diverse methodological branches stemming from MSP [12] and Energy [24] paradigms. LINe [2] innovates by reducing neuron-induced noise through the calculation of Shapley values. Yu et al. [44] distinguish ID from OOD data by identifying neural network blocks with optimal differentiation based on the norms of their features. DICE [34] improves discrimination by pruning weights in the fully connected layer according to the contribution units make during classification. On the other end of the spectrum, entirely computation-free post hoc methods such as ReAct [33] and ASH [8] have shown promise. ReAct [33] investigates activations prior to the fully connected layer, applying rectification to suppress extreme activations that OOD data tend to trigger, thereby achieving refined detection outcomes. Similarly, ASH [8] prunes the activations inputted to the fully connected layer, but it achieves even more enhanced results compared to DICE [34] by its selective pruning strategy. In this paper, our comparison mainly focuses on post hoc methods, since our method also belongs to this category.
@@ 3 Neural Activation Prior
Our NAP is based on the following observation for OOD detection: for a channel located before the global pooling layer in a fully trained neural network, the likelihood that a small number of its neurons activated with a stronger response to an ID sample is significantly higher compared to an OOD sample. For the behavior in other layers of the neural network, refer to the discussion around Figure 5 in Section 5.
To formally describe this observation, we first define the concept of neural activation. Consider a trained classification neural network $f$, assuming it receives $D$-dimensional input $x$ and outputs $K$-dimensional logits. That is $f:\mathbb{R}^{D}\rightarrow\mathbb{R}^{K}$. We concentrate on the activation tensor $\mathbf{A}$ , located at the penultimate layer just before the global pooling operation, as illustrated in Figure 2(a). Let the dimensions of $\mathbf{A}$ be $C\times H\times W$, where $C$ is the number of channels, and $H$ and $W$ are the spatial dimensions.
(a) NAP in CNN
(b) NAP in Transformer
Figure 3: Illustration identifying the focus zone of Neural Activation Prior (NAP) in classification neural networks. The figure highlights the specific position of NAP in the network: (a) For the CNN-based model, the activation value in the green part is the location of the NAP proposed in this article, which is before global pooling. (b) For the Transformer model, NAP is highlighted within the cls token’s attention weights in the last Transformer block, illustrating a targeted approach in contrast to most existing methods that focus on regions after global pooling.
@@
    where $\text{Max}(\mathbf{A}_{j})$ is the maximum value among all elements in the activation vector $\mathbf{A}_{j}$. Notice that $\mathbf{A}_{jkl}$ is processed by ReLU (as shown in Figure 2(a)), hence $\mathbf{A}_{jkl}\geq 0$.
@@
Inspired by the concept of signal-to-noise ratio, we interpret $\text{Max}(\mathbf{A}_{j})$ as the signal strength and $\text{Mean}(\mathbf{A}_{j})$ as the noise strength. The ratio of these values can be viewed as an indicator of the quality of the activation vector $\mathbf{A}_{j}$. Note that we calculate this metric separately for each channel in the network, since different channels are usually used to detect different patterns. In the context of OOD detection, this quality measure can be used to assess whether the neural network recognizes the input sample — in other words, this measure can be used to judge whether a sample is within the training set distribution. As shown in Figure 1, this phenomenon is termed as the neural activation prior.
It is worth noting that our proposed prior is orthogonal to existing OOD detection methods. As illustrated in Figure 2(a), existing methods mainly focus on the network output and weight of the penultimate layer after global pooling, and leverage them to design various scoring functions for OOD detection. In contrast, the prior we propose focuses on the channels of the penultimate layer before global pooling. Since the distribution information within these channels is inevitably lost during the global pooling process, our proposed prior is complementary to existing work.
Extension to Transformer Backbones. We observe that the classify (cls) token in the last block of Transformers can be effectively utilized as an analogue to the pooled activations used in CNNs for our method. Consequently, as illustrated in Figure 2(b), we calculate NAP score by employing the attention vectors associated with the cls token from the final Transformer block. This approach mirrors our methodological framework in CNNs, facilitating a coherent extension across both architectures. The specifics of the scoring function for NAP, applicable to both CNNs and Transformers, will be further elucidated in Section 4.2.